OpenStudy (anonymous):

When the equation 3x - 4y = 24 is written in slope-intercept form, what is the value of m?

OpenStudy (anonymous):

It's 3/4, correct?

OpenStudy (nubeer):

y = mx + c rearrange your given equation to make a format like this 3x-4y =24 3x - 24 = 4y now divide both sides by 4 .. what will u get?

OpenStudy (nubeer):

yes its correct.

OpenStudy (anonymous):

is the answer 3/4 or 9